Identify the relevant elasticity concept for the given event. Then, compute the measure of elasticity using average prices and quantities in your calculations. In all cases, assume that these are ceteris paribus changes.


As average household income in Canada increases by 10 percent, annual sales of Toyota Camrys increase from 56,000 to 67,000. The relevant elasticity concept is _________ and the computed elasticity is ___________


1
Expert's answer
2022-02-02T12:37:50-0500

Elasticity Concept

The measure of Toyota Camrys demand sensitivity with respect to the changes in Canada's household income.

Computed Elasticity

"elasticity =\frac{\% change\ in\ quantity\ demanded}{\% change\ in\ price}"

for this case, we shall replace the "\% change\ in\ price" with "\% change\ in\ income"

"\% change\ in\ Quantity\ demanded=\frac{11000}{56000}\times100=19.6"

"\%change\ in\ income=10\%"

therefore;

"\bold{Computed\ elasticity}=\frac{19.6}{10}=1.96"